OpenStudy (anonymous):

Complete the ordered pairs so they are solutions of the equation: -x+6y=-24

OpenStudy (anonymous):

I have 2 of them already. (24,0) (0,-4) the last one is (__,-1)

OpenStudy (anonymous):

(18,-1)

OpenStudy (anonymous):

since -18 - 6 = -24

OpenStudy (anonymous):

oh ok. thanks so much!!

OpenStudy (anonymous):

Can you help with another?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Complete the ordered pairs so they are solutions of the equation: 3x-6y=-18. I already have (-6,0), (0,3) and the last one is (__,3)

OpenStudy (anonymous):

(0,3)

OpenStudy (anonymous):

lol you had two with same y coordinates

OpenStudy (anonymous):

(-12,-3) works too

OpenStudy (anonymous):

wow. i did not even notice that. it was right though so thanks so much :) i tried -12 and it wasnt right.

OpenStudy (anonymous):

what about this one? 3x+2y=12? (4,0) (0,6) snd (__,3)